# Transverse vortices induced by modulated granular shear flows of elongated particles

### v1

@@ IV.1 Governing equations
Taking linear terms in $\epsilon$ in Eqs. 10 and 11, we have
@@
$\rho\phi_{c}\dot{\gamma}_{0}d\left(U_{j}\partial_{j}u_{0,i}+u_{0,j}\partial_{j}U_{i}\right)=$
@@
Substituting the previous expressions and using the fact that $\partial_{x}(.)=0$, we obtain:
@@
which, once rearranged and simplified (using in particular $\partial_{y}U_{y}+\partial_{z}U_{z}=0$), gives:
@@
With this, we finally obtain the following governing linear equations:
@@
IV.2 A single mode solution
Simple modes of the above equations (33-36) are the form of the product of sine and cosine functions:
@@ IV.2 A single mode solution
with dimensionless wavenumbers $k_{y}=2\pi d/L_{y}$ and $k_{z}=2\pi d/L_{z}$. Assuming the perturbation itself has a profile along a mode:
@@
then Eqs. (33-36) simplify into:
@@
where we have defined $\beta_{c}=\eta_{c}\nu_{2c}/\mu_{c}^{2}$ for short.
@@
as they should, so that there is no flux through the sides of the cell. Furthermore, the constant $-a_{p}$ can be added to $P$ to make it vanish at the cell corners. This solution shows a velocity field with four vortices in the $y-z$ plane, as displayed in Fig. 1c and Fig. 2a.
@@
Figure 2: Comparison of the DEM simulations (red) and calculations based on the generalized granular rheology (blue) for $\alpha{=}2$, $\mu_{p}{=}0.1$, and $L_{z}/L_{y}{=}1$ (presented also in Fig. 4b). (a): The vortex pattern of the $y{-}z$ plane obtained by the coarse-grained averaging of the velocities in simulations, and the expected pattern according to the theory. The width of the arrows represents the magnitude of velocities. (b): Cross-sections of the velocity profiles at $y{=}3.5$ (continuous lines) and $z{=}7.5$ (dashed lines). In the case of simulations, we present two cross-sections obtained by applying different perturbation strengths. Stronger perturbation corresponds to the light red (thicker lines, $A{=}2{\cdot}10^{-4}$), while weaker perturbation is represented by the dark red (narrower lines, $A{=}1{\cdot}10^{-4}$).
@@ V.1 Numerical setup
In order to study the behavior of a sheared granular system in the presence of spatially modulated additional body force we performed DEM simulations in the 3D plane Couette geometry (see Fig. 1), with periodic boundary conditions in the $x$ (flow) and $z$ (neutral) directions and Lees-Edwards boundary conditions in the $y$ (velocity gradient) direction. This way wall-effects could be eliminated. The particles were frictional sphero-cylinders with length-to-diameter aspect ratio $\alpha=\ell/d$. In addition, to avoid crystallization, some amount of size polydispersity was introduced: while we kept the aspect ratio constant, we drew the diameters of the particles from a uniform distribution with a ratio of standard deviation to mean of $10\%$. The length, time, and mass units of the simulation were set implicitly by setting the mean particle diameter $d$, density $\rho$, and contact stiffness $k$ (equal for the normal and tangential force) to unity.
We created the initial conditions of random particle orientation with overdamped dynamics and afterward sheared the system at constant shear rate $\dot{\gamma}$. During shear, we employed stress control, where one side of the box, $L_{y}$, was adjusted by a feedback loop such that the corresponding normal stress $\sigma_{yy}$ fluctuated around a specified value $-p_{y}=10^{-3}$. We kept $L_{x}$ and $L_{z}$ fixed. For more details about the DEM simulations, such as the interaction forces, we refer the reader to Refs. Nagy et al. (2017, 2020).
To implement the external body perturbation (term $b_{i}$ in RHS of Eq. 11), we applied a position-dependent force at the center of every particle, acting in the $x$-direction:
@@
The constant $A$ is typically set to a small value in the order of $10^{-4}$ so that it acts as a perturbation with respect to the main shear flow – see next section for comments about this point. It relates to $\epsilon$ introduced in the previous section as $A=\epsilon\frac{\Pi_{0}}{d}V_{P}/\phi_{c}$, where $V_{P}$ is the average volume of the particles.
We ran several DEM simulations to measure the response of the system and compare it to the theoretical predictions. We varied the particle aspect ratio ($\alpha$), the particle-particle friction coefficient ($\mu_{p}$), the strength of the perturbation ($A$), and the cell aspect ratio ($L_{z}/L_{y}$). For each case, we ran $5$ independent realizations and averaged over them. All measurements were taken in the stationary state after a deformation of at least $\gamma=10$. All quantities of interest were time-averaged for a duration corresponding to a deformation of at least $\gamma=100$. The number of particles in the cell was $4000$, except for the largest cell aspect ratio for which it was increased to $10300$.
@@ V.2 Calculation of the response
To obtain the response of the system to the force perturbation, we calculated the coarse-grained (CG) Goldhirsch (2010) velocity profiles in the $y-z$ plane using the particle length $\ell$ as the CG scale and also averaging in the $x$-direction along which the system is homogeneous. This resulted in the velocity vector $\bm{u}$ (see Eq. 13). After subtracting the base velocity $\bm{u_{0}}$ and dividing by $\epsilon\dot{\gamma}_{0}d$, we obtained the rescaled velocity response $U_{i}(y,z)$ (here for $i=y,z$ only), which have been fitted according to Eqs. 37-39. This way we obtain the values of the coefficients $a_{i}$, which can be directly compared to those predicted by the theory in Eqs. 46-49. As said above, to compute the value of $\epsilon$ from that of $A$, one needs the main pressure $\Pi_{0}$, which was calculated from the values of $N_{1}$, $N_{2}$ (taken from Nagy et. al Nagy et al. (2020) in the limit $I\to 0$, see also Fig. A1 in Appendix A), and the imposed normal stress $\sigma_{yy}$. As $N_{1}$ and $N_{2}$ depend on $\alpha$ and $\mu_{p}$, the value of $\Pi_{0}$ varied from case to case but was typically in the order of $\Pi_{0}\sim 10^{-3}$, i.e. expectedly comparable to $\sigma_{yy}$.
@@ VI Results and discussion
First, we present cross sections of the velocity field in the $y-z$ plane, which show the secondary circulation resulting from the modulated forcing in the $x$-direction in a sheared system (as sketched in Fig. 1). In Fig. 2 we show the results of simulations and theory for a cell with aspect ratio $L_{z}/L_{y}=1$, filled with particles of aspect ratio $\alpha=2$ and friction coefficient $\mu_{p}=0.1$. The overall agreement between the numerics (red) and the calculations (blue) made according to the generalized granular rheology is good, they both show two pairs of counterrotating vortices. In panel (a) we show the CG velocity field in the $y{-}z$ plane, while panel (b) compares cross-sections of the velocity field at $y=3.5$ (continuous line) and $z=7.5$ (dashed line), for the $y$ and $z$-components of velocity, and one can see the sinusoidal shape in both cases, as anticipated by equations 38-39.
@@
We have tested different perturbation strengths $A=1\cdot 10^{-4}$, $2\cdot 10^{-4}$, $3\cdot 10^{-4}$, and $5\cdot 10^{-4}$. This perturbation has to be strong enough to allow for a measure of the response out of the noise, but ideally not too large to be compatible with the linear analysis developed above. Such values $A\simeq 10^{-4}$ correspond to $\epsilon$ on the order of a few percents, and this is indeed the relative magnitude of the transverse components of the velocity response. However, for the same $A$, that of the stream-wise velocity is significantly larger, on the order of a few tens of percents. This can actually be understood from the linear theory: comparing the response coefficients $a_{x}$ and $a_{y}$ in Eqs. 46 and 47 respectively, one can see that $a_{x}$ must be at least larger than $a_{y}$ by a factor of $\mu_{c}/\nu_{2c}\simeq 10$. Besides, as the numerator of $a_{x}$ comes as a sum, whereas that of $a_{y}$ as a difference, this provides an additional factor of a few units. As a consequence, the response to the perturbation in the $x$-direction is significantly larger than that in the transverse direction. This inevitably puts us away from a proper linear regime as we cannot reliably measure the transverse response with $A$ smaller than $10^{-4}$ (too much noise for reasonable time averaging – a deformation $\gamma\simeq 200$). However, we observe that, for the two smallest values of the perturbation we have used ($A=1\cdot 10^{-4}$ and $2\cdot 10^{-4}$), the transverse velocity response is linear in the sense that, not only the profile shape is sinusoidal, but the curves collapse once rescaled by $A$, as shown in Fig. 2b. For the larger values of $A$ we have tested, the strength of the response can not be rescaled anymore.
@@
A strong and robust prediction of the theory is that the sense of rotation of the vortices should change when the cell aspect ratio $L_{z}/L_{y}$ increases. Indeed, our numerical simulations show this inversion as we see in the cross sections of the velocity fields presented in Fig. 3a, or in Fig.4 where we see the change of sign of the vortex amplitude as a function of $L_{z}/L_{y}$. The origin of this inversion in the theoretical description can be seen in the expressions for the amplitudes $a_{y}$ and $a_{z}$ (47-48), where the numerator changes sign for $k_{y}/k_{z}=L_{z}/L_{y}=\sqrt{\pi_{c}/(\pi_{c}+\nu_{2c})}\gtrsim 1$, while the denominator stays positive.
Figure 3: (a) Coarse-grained velocity profiles of cells with different cell aspect ratios obtained by the DEM simulations for $\alpha{=}1.5$, $\mu_{p}{=}0.01$, and $A{=}2{\cdot}10^{-4}$. The width of the arrows represents the magnitude of velocities. For the time evolution of the vortex pattern, we refer the reader to the SM videos SM0 (a, b). Corresponding packing fraction (b) and velocity gradient $|\text{grad}(u_{x})|$ (c) maps. Note that for the largest cell aspect ratio $L_{z}/L_{y}{=}3.18$, the particle number was increased to $10300$, while in all the other cases it is $4000$. Note also that the colorbars of panels (b) and (c) are reversed to emphasize the similarity of the patterns.
Our numerical simulations also allow us to extract information about the spatial distribution of the packing fraction in the system. In Figs. 3b-c we show the maps for the packing fraction as well as for the norm of the gradient of the $x$-component of the velocity $|\text{grad}(u_{x})|$ corresponding to the systems presented in Figs. 3a. As we see, there is a correlation between shear rate and packing fraction which can be related to two effects: (i) In our system the inertial number $I$ reaches a value of $10^{-2}$, meaning that inertial effects are already present, although with small amplitude. Such effects lead to slightly decreased density due to more intensive collisions at the locations with larger shear rates. (ii) Spatial variation of the direction of the shear gradient will lead to a spatial modulation of the shear-induced alignment of the grains which can also lead to decreased packing fraction if it causes frustration in the grain orientation. Irrespective of the cause of the spatial variation of the packing fraction, the most important aspect is that we see that for the secondary flow, the change in the rotation direction is not accompanied by a change in the packing fraction map.
In Fig. 4, we present data for several coefficients of friction, and for two values of the particle aspect ratio, $\alpha=1.5$ and $2$. For each case, we prepared simulations with four different cell aspect ratios. The symbols show the results of the DEM simulations, while the theoretical prediction is represented by the dashed lines. To present the response of the system, we use a rescaled value, namely $a_{y\text{,rescaled}}=a_{y}(L_{y}^{2}+L_{z}^{2})/(L_{y}^{2}L_{z}^{2})$, as suggested by the pre-factor in expressions  47-48. This way, the role of the cell aspect ratio is highlighted, and the effect of the actual dimensions of the cell is ruled out, facilitating the comparison between the different systems. One can see the good qualitative agreement between the simulated and calculated data in terms of changes in the vortex amplitude as a function of $L_{z}/L_{y}$. In addition, the inversion can also be tracked in Appendix A in Figs. A1 and A2 where the quantities $\mu$, $\nu_{2}$, $\pi$ and $a_{y}$ are presented as a function of $\mu_{p}$ and $\alpha$.
Further investigating the data we see that the $a_{y}$ vs. $L_{z}/L_{y}$ curves are not independent of microscopic parameters. For example in Fig. 4a, for the numerical data the zero-crossing is observed at larger cell aspect ratio for smaller $\mu_{p}$. This means that for an intermediate but fixed $L_{z}/L_{y}$, the sign of $a_{y}$ can be changed when varying the particle friction coefficient. The theoretical lines display similar behavior, but for a much more restricted range of $L_{z}/L_{y}$, and since the lines cross the other way round, the effect is opposite (see Appendix A Fig. A2a). Finally, note that each point obtained by DEM simulations is an average of $5$ independent realizations, but the difference between these realizations is small, and we have not seen a change of sign between the different realizations corresponding to the same simulation parameters.
@@
Figure 4: Response of the system, rescaled by the system size ($a_{y\text{,rescaled}}{=}a_{y}(L_{y}^{2}{+}L_{z}^{2})/(L_{y}^{2}L_{z}^{2})$), as a function of cell aspect ratio for two particle aspect ratios $\alpha{=}1.5$ (a) and $\alpha{=}2$ (b), and for different particle coefficients of friction $\mu_{p}$ (legends). The strength of perturbation is $A{=}2{\cdot}10^{-4}$. Dashed lines show the expected values according to the theory. The symbols show the results of simulations, averaged over $5$ different realizations. Their span is represented by the light-shaded background of the lines. Importantly, note the different vertical axes on these graphs for simulation results (left) and theoretical predictions (right) – a qualitative-only agreement is reached.
@@
We note that the agreement between simulations and theory is qualitative only: the theory predicts a larger amplitude for the response compared to what we get from numerics. This is why we have used distinct vertical axes in the figures for the magnitude of the velocities. We also see that the difference between the simulation results and the theoretical prediction is larger for large $L_{z}/L_{y}$ than for small $L_{z}/L_{y}$, and that the inversion is observed at a larger $L_{z}/L_{y}$ than predicted. The lack of quantitative matching could be thought to be surprising, given the fact that we are using calibrated parameters $\mu_{c}$, $\nu_{1c}$ and $\nu_{2c}$ of the rheological law on very similar numerical runs (in the homogeneous case, i.e. with $F_{x}=0$) Nagy et al. (2020). The possible adjustment of the remaining free parameter $\eta_{c}$ does not improve much the matching, as the theoretical results are in practice not very sensitive to its value – and all presented curves have been computed with $\eta_{c}$ (or $\beta_{c}$) equals 00. Several reasons can be invoked to explain this quantitative difference. First, as mentioned in the previous section, even for a perturbation amplitude as small as possible to measure a transverse velocity response on the order of a few percent, the corresponding response in the $x$-direction is large, significantly above the linear regime. Also, one must recall that the theory is for continuous fields while the simulations are performed with discrete grains. This leads to the question of whether or not the ratio between the system and the grain sizes is large enough. Although it seems fine for the base (homogeneous) system, it is not obvious that when heterogeneities or gradients are introduced, such as those induced by the forcing $F_{x}$, they are sufficiently resolved with the total number of grains we have in the runs. These spatial gradients are also known to induce non-local effects Pouliquen and Forterre (2009); Bouzid et al. (2013); Henann and Kamrin (2013), which are not accounted for in the present theory. A good test would be of course to run simulations with a larger number of grains, especially in the case of more elongated particles. This is however computationally expensive, especially as averaging needs to be performed over rather long deformations. A last reason to explain the predicted vs. measured quantitative mismatch is that the theory assumes incompressibility. However, as the packing fraction maps show (Fig. 3b), this is not the case in the DEM simulations. Clearly, lower packing fractions correspond to higher velocity gradients (Fig. 3c). These heterogeneities are likely to result from the strong stream-wise perturbation and can have significant feedback on the secondary flows and vortex amplitude, in a way that needs to be further investigated, but which is beyond the scope of this paper.
@@
Despite this quantitative mismatch between theory and simulations, evidence for the inversion of the vortex pattern and its geometrical origin is a major result of the present paper. Also, following the rheological interpretation of the numerical measurements, we understand that these secondary flows are induced by the normal stress differences, and in particular by the second one, as the factor $\nu_{2c}$ is present in the expressions of the vortex amplitudes $a_{y}$ and $a_{z}$ (see Eqs. 47-48). Interestingly, a buoyancy-driven explanation for secondary flows in Couette cells was proposed by Krishnaraj and Nott (2016). Here, although we do observe a packing fraction pattern associated with the vortices (Fig. 3b), it cannot generate convection as there is no gravity in the system. Besides, we do not observe any qualitative change in these packing fraction maps, nor for those of the velocity gradient (Fig. 3c), when the cell aspect ratio increases, associated with the vortex pattern inversion.

### v2

@@ IV.1 Governing equations
Taking linear terms in $\epsilon$ in Eqs. (10) and (11), we have
@@
$\rho\phi_{c}\dot{\gamma}_{0}d\left(U_{j}\partial_{j}u_{0,i}+u_{0,j}\partial_{j}U_{i}\right)$
@@
Substituting the previous expressions and using the fact that $\partial_{x}(.)=0$, we obtain
@@
which, once rearranged and simplified (using in particular $\partial_{y}U_{y}+\partial_{z}U_{z}=0$), gives
@@
With this, we finally obtain the following governing linear equations
@@
To recall, the first of these four equations comes from mass conservation assuming incompressibility Eq. (10) and the next three express momentum conservation Eq. (11).
IV.2 Single-mode solution
Simple modes of the above Eqs. (33) to (36) are the form of the product of sine and cosine functions:
@@ IV.2 Single-mode solution
with dimensionless wavenumbers $k_{y}=2\pi d/L_{y}$ and $k_{z}=2\pi d/L_{z}$. Assuming the perturbation itself has a profile along a mode
@@
then Eqs. (33) to (36) simplify into
@@
where we define $\beta_{c}=\eta_{c}\nu_{2c}/\mu_{c}^{2}$ for short. Note that $a_{y}$ and $a_{z}$ are simply proportional, as expected from a velocity field of vanishing divergence, but homogeneous in the $x$ direction Eq. (24): $a_{z}=-a_{y}\,k_{y}/k_{z}$.
@@
as they should, so that there is no flux through the sides of the cell. Furthermore, the constant $-a_{p}$ can be added to $P$ to make it vanish at the cell corners. This solution shows a velocity field with four vortices in the $y$-$z$ plane, as displayed in Figs. 1(c) and 2(a).
@@ V.1 Numerical setup
To study the behavior of a sheared granular system in the presence of spatially modulated additional body force we performed DEM simulations in the three-dimensional plane Couette geometry (see Fig. 1), with periodic boundary conditions in the $x$ (flow) and $z$ (neutral) directions and Lees-Edwards boundary conditions in the $y$ (velocity gradient) direction. This way wall effects could be eliminated. The particles were frictional spherocylinders with length-to-diameter aspect ratio $\alpha=\ell/d$. In addition, to avoid crystallization, some amount of size polydispersity was introduced: while we kept the aspect ratio constant, we drew the diameters of the particles from a uniform distribution with a ratio of standard deviation to mean of $10\%$. The length, time, and mass units of the simulation were set implicitly by setting the mean particle diameter $d$, density $\rho$, and contact stiffness $k$ (equal for the normal and tangential force) to unity.
We created the initial conditions of random particle orientation with overdamped dynamics, and afterward sheared the system at constant shear rate $\dot{\gamma}$. During shear, we employed stress control, where one side of the box, $L_{y}$, was adjusted by a feedback loop such that the corresponding normal stress $\sigma_{yy}$ fluctuated around a specified value $-p_{y}=10^{-3}$. We kept $L_{x}$ and $L_{z}$ fixed. The simulations were run with an inertial number on the order of $I\simeq 10^{-2}$, which is a good compromise between a shear rate small enough for the system to behave in the quasistatic regime, but large enough to achieve significant deformation in a reasonable CPU time. Recall that deformations as large as $\gamma\simeq 200$ are required to see secondary flows that average out from the noise. For more details about the DEM simulations, such as the interaction forces, we refer the reader to Nagy et al. (2017, 2020).
To implement the external body perturbation [term $b_{i}$ in RHS of Eq. (11)], we applied a position-dependent force at the center of every particle, acting in the $x$ direction:
@@
The constant $A$ is typically set to a small value in the order of $10^{-4}$ so that it acts as a perturbation with respect to the main shear flow. It relates to $\epsilon$ introduced in the previous section as $A=\epsilon\frac{\Pi_{0}}{d}V_{P}/\phi_{c}$, where $V_{P}$ is the average volume of the particles. Importantly, as discussed in detail in the next section, this small value for $A$ already induces significant nonlinearities for the streamwise response, although lower values would make the measurement of the transverse response (the vortices we are interested in) very difficult given the cell deformation we can reach ($\gamma\simeq 200$) and the noise associated with grain collisions.
We ran several DEM simulations to measure the response of the system and compare it to the theoretical predictions. We varied the particle aspect ratio ($\alpha$), the particle-particle friction coefficient ($\mu_{p}$), the strength of the perturbation ($A$), and the cell aspect ratio ($L_{z}/L_{y}$). For each case, we ran five independent realizations and averaged over them. All measurements were taken in the stationary state after a deformation of at least $\gamma=10$. All quantities of interest were time-averaged for a duration corresponding to a deformation of at least $\gamma=100$. The number of particles in the cell was $4000$, except for the largest cell aspect ratio for which it was increased to $10300$.
@@ V.2 Calculation of the response
To obtain the response of the system to the force perturbation, we calculated the coarse-grained (CG) Goldhirsch (2010) velocity profiles in the $y$-$z$ plane using the particle length $\ell$ as the CG scale and also averaging in the $x$ direction along which the system is homogeneous. This resulted in the velocity vector $\bm{u}$ [see Eq. (13)]. After subtracting the base velocity $\bm{u_{0}}$ and dividing by $\epsilon\dot{\gamma}_{0}d$, we obtained the rescaled velocity response $U_{i}(y,z)$ (here for $i=y,z$ only), which have been fitted according to Eqs. (37) to (39). This way we obtain the values of the coefficients $a_{i}$, which can be directly compared to those predicted by the theory in Eqs. (46) to (49). As said above, to compute the value of $\epsilon$ from that of $A$, one needs the main pressure $\Pi_{0}$, which was calculated from the values of $N_{1}$, $N_{2}$ (taken from Nagy et al Nagy et al. (2020) at the value of $I=10^{-2}$, see also Fig. 5 in Appendix A), and the imposed normal stress $\sigma_{yy}$. As $N_{1}$ and $N_{2}$ depend on $\alpha$ and $\mu_{p}$, the value of $\Pi_{0}$ varied from case to case but was typically in the order of $\Pi_{0}\sim 10^{-3}$, i.e., expectedly comparable to $\sigma_{yy}$.
Figure 3: (a) Coarse-grained velocity profiles of cells with different cell aspect ratios obtained by the DEM simulations for $\alpha=1.5$, $\mu_{p}=0.01$, and $A=2\times 10^{-4}$. The width of the arrows represents the magnitude of velocities. For the time evolution of the vortex pattern, we refer the reader to the SM videos SM . (b) Corresponding packing fraction and (c) velocity gradient $|\text{grad}(u_{x})|$ maps. Note that for the largest cell aspect ratio $L_{z}/L_{y}=3.18$, the particle number was increased to $10300$, while in all the other cases it is $4000$. Note also that the colorbars of panels (b) and (c) are reversed to emphasize the similarity of the patterns.
@@ VI Results and discussion
First, we present cross sections of the velocity field in the $y$-$z$ plane, which show the secondary circulation resulting from the modulated forcing in the $x$ direction in a sheared system (as sketched in Fig. 1). In Fig. 2 we show the results of simulations and theory for a cell with aspect ratio $L_{z}/L_{y}=1$, filled with particles of aspect ratio $\alpha=2$ and friction coefficient $\mu_{p}=0.1$. The overall agreement between the numerics (red) and the calculations (blue) made according to the generalized granular rheology is good; they both show two pairs of counterrotating vortices. In Fig. 2(a) we show the CG velocity field in the $y$-$z$ plane, while Fig. 2(b) compares cross sections of the velocity field at $y=3.5$ (continuous line) and $z=7.5$ (dashed line), for the $y$ and $z$ components of velocity, and one can see the sinusoidal shape in both cases, as anticipated by Eqs. (38) and (39).
@@
We tested different perturbation strengths $A=1\times 10^{-4}$, $2\times 10^{-4}$, $3\times 10^{-4}$, and $5\times 10^{-4}$. This perturbation has to be strong enough to allow for a measure of the response out of the noise, but ideally not too large to be compatible with the linear analysis developed above. Such values $A\simeq 10^{-4}$ correspond to $\epsilon$ on the order of a few percents, and this is indeed the relative magnitude of the transverse components of the velocity response. However, for the same $A$, that of the stream-wise velocity is significantly larger, on the order of a few tens of percents. This can actually be understood from the linear theory: comparing the response coefficients $a_{x}$ and $a_{y}$ in Eqs. (46) and (47) respectively, one can see that $a_{x}$ must be at least larger than $a_{y}$ by a factor of $\mu_{c}/\nu_{2c}\simeq 10$. In addition, as the numerator of $a_{x}$ comes as a sum, whereas that of $a_{y}$ comes as a difference, this provides an additional factor of a few units. As a consequence, the response to the perturbation in the $x$ direction is significantly larger than that in the transverse direction. This inevitably puts us away from a proper linear regime as we cannot reliably measure the transverse response with $A$ smaller than $10^{-4}$ (too much noise for reasonable time averaging, a deformation $\gamma\simeq 200$). However, we observe that, for the two smallest values of the perturbation we have used ($A=1\times 10^{-4}$ and $2\times 10^{-4}$), the transverse velocity response is linear in the sense that, not only the profile shape is sinusoidal, but the curves collapse once rescaled by $A$, as shown in Fig. 2(b). For the larger values of $A$ we have tested, the strength of the response can not be rescaled anymore.
@@
A strong and robust prediction of the theory is that the sense of rotation of the vortices should change when the cell aspect ratio $L_{z}/L_{y}$ increases. Indeed, our numerical simulations show this inversion as we see in the cross sections of the velocity fields presented in Fig. 3(a), or in Fig. 4 where we see the change of sign of the vortex amplitude as a function of $L_{z}/L_{y}$. The origin of this inversion in the theoretical description can be seen in the expressions for the amplitudes $a_{y}$ and $a_{z}$ in Eqs. (47) and (48), where the numerator changes sign for $k_{y}/k_{z}=L_{z}/L_{y}=\sqrt{\pi_{c}/(\pi_{c}+\nu_{2c})}\gtrsim 1$, while the denominator stays positive.
@@
Figure 4: Response of the system, rescaled by the system size [$a_{y\text{,rescaled}}{=}a_{y}(L_{y}^{2}{+}L_{z}^{2})/(L_{y}^{2}L_{z}^{2})$], as a function of cell aspect ratio for two particle aspect ratios (a) $\alpha{=}1.5$ and (b) $\alpha{=}2$, and for different particle coefficients of friction $\mu_{p}$ (legends). The strength of perturbation is $A=2\times 10^{-4}$. Dashed lines show the expected values according to the theory. The symbols show the results of simulations, averaged over five different realizations. Their span is in the order of the point size. Importantly, note the different vertical axes on these graphs for simulation results (left) and theoretical predictions (right) — a qualitative-only agreement is reached.
Our numerical simulations also allow us to extract information about the spatial distribution of the packing fraction in the system. In Figs. 3(b) and 3(c), we show the maps for the packing fraction as well as for the norm of the gradient of the $x$ component of the velocity $|\text{grad}(u_{x})|$ corresponding to the systems presented in Figs. 3(a). As we see, there is a correlation between shear rate and packing fraction, which can be related to two effects (i) In our system, the inertial number $I$ reaches a value of $10^{-2}$, meaning that inertial effects are already present, although with small amplitude. Such effects lead to slightly decreased density due to more intensive collisions at the locations with larger shear rates. (ii) Spatial variation of the direction of the shear gradient will lead to a spatial modulation of the shear-induced alignment of the grains, which can also lead to decreased packing fraction if it causes frustration in the grain orientation. Irrespective of the cause of the spatial variation of the packing fraction, the most important aspect is that we see that for the secondary flow, the change in the rotation direction is not accompanied by a change in the packing fraction map.
In Fig. 4, we present data for several coefficients of friction, and for two values of the particle aspect ratio, $\alpha=1.5$ and $2$. For each case, we prepared simulations with four different cell aspect ratios. The symbols show the results of the DEM simulations, while the theoretical prediction is represented by the dashed lines. To present the response of the system, we use a rescaled value, namely $a_{y\text{,rescaled}}=a_{y}(L_{y}^{2}+L_{z}^{2})/(L_{y}^{2}L_{z}^{2})$, as suggested by the prefactor in expressions (47) and (48). This way, the role of the cell aspect ratio is highlighted, and the effect of the actual dimensions of the cell is ruled out, facilitating the comparison between the different systems. One can see the good qualitative agreement between the simulated and calculated data in terms of changes in the vortex amplitude as a function of $L_{z}/L_{y}$. In addition, the inversion can also be tracked in Appendix A in Figs. 5 and 6 where the quantities $\mu$, $\nu_{2}$, $\pi$ and $a_{y}$ are presented as a function of $\mu_{p}$ and $\alpha$.
Further investigating the data, we see that the $a_{y}$ versus $L_{z}/L_{y}$ curves are not independent of microscopic parameters. For example, in Fig. 4a, for the numerical data, the zero-crossing is observed at a larger cell aspect ratio for smaller $\mu_{p}$. This means that for an intermediate but fixed $L_{z}/L_{y}$, the sign of $a_{y}$ can be changed when varying the particle friction coefficient. The theoretical lines display similar behavior, but for a much more restricted range of $L_{z}/L_{y}$, and since the lines cross the other way round, the effect is opposite [see Appendix A Fig. 6(a)]. Finally, note that each point obtained by DEM simulations is an average of five independent realizations, but the difference between these realizations is small, and we have not seen a change of sign between the different realizations corresponding to the same simulation parameters. For comparison, we present the $a_{x}$ versus $L_{z}/L_{y}$ curves in Fig. 7 of Appendix A.
Figure 5: (a), (d) $\mu$, (b), (e) $\nu_{2}$, and (c), (f) $\pi$ as a function of the particle-particle coefficient of friction (a)-(c) ($\mu_{p}$) and (d)-(f) the particle aspect ratio ($\alpha$) for cell aspect ratio $L_{z}/L_{y}=1$ and inertial number $I=10^{-2}$ using results of Nagy et al. Nagy et al. (2020).
@@
We note that the agreement between simulations and theory is qualitative only: the response measured in the numerical simulations is typically of smaller amplitude than what the theory always predicts. This is why we used distinct vertical axes in the figures for the magnitude of the velocities. We also see that the difference between the simulation results and the theoretical prediction for the transverse direction is larger for large $L_{z}/L_{y}$ (order of 20) than for small $L_{z}/L_{y}$ (order of two to three), and that the inversion point, i.e. the zero-crossing of the curves in Fig. 4, is observed at a larger $L_{z}/L_{y}$ than predicted. As for the response in the shear direction (Fig. 7), the difference between the theory and numerics does not depend on the cell aspect ratio, it is in the order of 20. The lack of quantitative matching could be thought to be surprising, given the fact that we are using calibrated parameters $\mu_{c}$, $\nu_{1c}$ and $\nu_{2c}$ of the rheological law on very similar numerical runs (in the homogeneous case, i.e. with $F_{x}=0$Nagy et al. (2020). The possible adjustment of the remaining free parameter $\eta_{c}$ does not improve much the matching, as the theoretical results are in practice not very sensitive to its value, and all presented curves have been computed with $\eta_{c}$ (or $\beta_{c}$) equals 00. We do not think that the rheological parameters of Nagy et al. (2020) should be reassessed. In fact, several reasons can be invoked to explain this quantitative difference. First, as mentioned in the previous section, even for a perturbation amplitude as small as possible to measure a transverse velocity response on the order of a few percent, the corresponding response in the $x$ direction is large, on the order a several tens of percent (see Fig. 7 in Appendix A), i.e. significantly above the linear regime. These nonlinearities are likely to be the main cause of the quantitative mismatch between simulations and theory. A second reason is that the theory is for continuous fields, while the simulations are performed with discrete grains. This leads to the question of whether or not the ratio between the system and the grain sizes is large enough. For the base homogeneous system, the transverse velocity gradient takes place over $\simeq 20$ grains, which seems to be reasonable. However, the heterogeneities we induce by the forcing $F_{x}$ operate at a scale typically four times smaller, i.e., over $\simeq 5$ grain diameters. Clearly, we are here at the limit of a continuum. Strong spatial gradients are also known to induce nonlocal effects Pouliquen and Forterre (2009); Bouzid et al. (2013); Henann and Kamrin (2013), which are not accounted for in the present theory. A good test would be, of course, to run simulations with a larger number of grains, especially in the case of more elongated particles. This is, however, computationally expensive, especially as averaging needs to be performed over rather long deformations. An extension of the present theory to account for the first nonlinear and/or nonlocal terms is a piece of work in itself, which we do not address here and leave for a future study. A last reason to explain the predicted versus measured quantitative difference is that the theory assumes incompressibility. However, as the packing fraction maps show [Fig. 3(b)], this is not the case in the DEM simulations. Clearly, lower packing fractions correspond to higher velocity gradients [Fig. 3(c)]. These heterogeneities are likely to result from the strong stream-wise perturbation and can have significant feedback on the secondary flows and vortex amplitude, in a nonobvious way that needs to be further investigated, but which is beyond the scope of this paper. Note finally that accounting for incompressible $\mu(I)$ rheology for granular flows requires advanced techniques such as those described in Barker et al. (2017); Barker and Gray (2017), leading to a substantial reformulation of the theory.
@@
Despite this quantitative mismatch between theory and simulations, evidence for the inversion of the vortex pattern and its geometrical origin is a major result of the present paper. Also, following the rheological interpretation of the numerical measurements, we understand that these secondary flows are induced by the normal stress differences, and in particular by the second one, as the factor $\nu_{2c}$ is present in the expressions of the vortex amplitudes $a_{y}$ and $a_{z}$ [see Eqs. (47) and (48)]. Interestingly, a buoyancy-driven explanation for secondary flows in Couette cells was proposed by Krishnaraj and Nott (2016). Here, although we do observe a packing fraction pattern associated with the vortices [Fig. 3(b)], it cannot generate convection as there is no gravity in the system. In addition, we do not observe any qualitative change in these packing fraction maps, nor for those of the velocity gradient [Fig. 3(c)], when the cell aspect ratio increases, associated with the vortex pattern inversion.
Figure 6: The rescaled response of the system [$a_{y\text{,rescaled}}=a_{y}(L_{y}^{2}{+}L_{z}^{2})/(L_{y}^{2}L_{z}^{2})$] as function of (a) the particle-particle coefficient of friction ($\mu_{p}$) and (b) the particle aspect ratio ($\alpha$) for cell aspect ratio $L_{z}/L_{y}=1$ (continuous lines), and $I=10^{-2}$. In panel (a) for $\alpha=1.5$, we show the influence of a slight change (±$\pm 10$%) of the cell aspect ratio on the response of the system (dashed green lines). For cell aspect ratios slightly larger than $1$, we have a change of sign in the response.